Opened 10 years ago

Closed 9 years ago

#228 closed flight processing (fixed)

GB08/08, flight day 068/2009, Wytham Woods

Reported by: benj Owned by: knpa
Priority: alpha 4 high Milestone: 2009 Data processing completion
Component: Archiving Keywords:
Cc: Other processors:

Description (last modified by knpa)

Data location: ~arsf/arsf_data/2009/flight_data/uk/GB08_08-2009_068_Wytham_Woods

Data arrived from ARSF via SATA J on 19th March 2009.

Scientific details: See /users/rsg/arsf/arsf_data/2008/ARSF_Applications/GB_and_NL/GB08-08 Proposal and Science Case.pdf

Priority: a4H

PI: R. Harding

Note - on SATA disk as GB08-05, but is actually 08-08

Sensors:

  • Eagle (27/04/2009)
  • Hawk (27/04/2009)
  • ALS-II Lidar (01/12/2009)
  • RCD (05/01/2010)

Quicklook

TBD.

Change History (29)

comment:1 Changed 10 years ago by benj

  • Description modified (diff)

comment:2 Changed 10 years ago by benj

  • Description modified (diff)

comment:3 Changed 10 years ago by benj

  • Description modified (diff)
  • Priority changed from medium to high

comment:4 Changed 10 years ago by mark1

  • Owner set to mark1

comment:5 Changed 10 years ago by mark1

GPT Timing offsets for E/H data:

Eagle line 2: -0.05

Eagle line 3: -0.01

Eagle line 4: -0.02

Hawk line 1: -0.05

Hawk line 2: -0.02

Hawk line 4: -0.02

comment:6 Changed 10 years ago by mark1

Due to issues with Eagle/Hawk sensor misalignment (see #220), have performed per line boresighting. Boresight values used to process data: (P,R,H)

Eagle line1: -0.22 -0.648 0.041

Eagle line 2: -0.077 -0.486 0.303

Eagle line 3: -0.234 -0.703 0.158

Eagle line 4: -0.197 -0.438 0.503

Hawk line 1: -0.1 0.261 0.05

Hawk line 2: -0.20 0.235 0.392

Hawk line 3: -0.311 0.211 0.245

Hawk line 4: -0.25 0.236 0.4

comment:7 Changed 10 years ago by mark1

  • Owner changed from mark1 to jatt

Delivery directory created at:

~mark1/scratch_space/2009_data/GB08_08-2009_068_Wytham_Woods/delivery/20090417/GB08_08/

But still needs new data quality report before sending.

comment:8 Changed 10 years ago by mark1

up to date data quality report added

comment:9 Changed 10 years ago by jatt

all looks good, rsynced to: /users/rsg/arsf/arsf_data/in_progress/2009/flight_data/uk/GB08_08-2009_068_Wytham_Woods

ready for delivery

comment:10 Changed 10 years ago by jatt

  • Owner changed from jatt to mark1

Eagle and hawk delivered 27/04/2009

comment:11 Changed 10 years ago by mggr

  • Priority changed from alpha 5 to alpha 4 high

Corrected grading following fixing trac's priority list..

comment:12 Changed 9 years ago by mark1

  • Owner mark1 deleted

comment:13 Changed 9 years ago by emca

  • Owner set to emca

Starting LIDAR processing

comment:14 Changed 9 years ago by emca

LIDAR processing is complete and project is now ready for delivery checking.
Delivery directory created at
~airborne/workspace/GB08_08-2009_068_Wytham_Woods/delivery

comment:15 Changed 9 years ago by mark1

  • Owner changed from emca to mark1
  • Status changed from new to assigned

comment:16 Changed 9 years ago by mark1

  • Owner changed from mark1 to emca

Delivery checked and looks good. Viewed data in GIS and lines up well. Very minor renaming edits made. Ready to deliver.

Data has not been rsynced back yet.

comment:17 Changed 9 years ago by emca

Lidar data dispatched to R. Harding on DVD 01/12/2009

comment:18 Changed 9 years ago by emca

LIDAR data has been rsynced to
/users/rsg/arsf/arsf_data/2009/flight_data/uk/GB08_08-2009_068_Wytham_Woods

comment:19 Changed 9 years ago by mark1

Photography delivered on hard disk 05/01/2010

comment:20 Changed 9 years ago by mark1

Photography delivery rsynced to

/users/rsg/arsf/arsf_data/2009/flight_data/uk/GB08_08-2009_068_Wytham_Woods

comment:21 Changed 9 years ago by emca

  • Component changed from Processing: general to Archiving
  • Owner changed from emca to benj

comment:22 Changed 9 years ago by knpa

  • Description modified (diff)

Started Archive Process

comment:23 Changed 9 years ago by knpa

  • Owner changed from benj to knpa

comment:24 Changed 9 years ago by knpa

No webcam files present

comment:25 Changed 9 years ago by knpa

Request made to DTG for archiving

comment:26 Changed 9 years ago by knpa

EDIT: recalled request as already archived by DTG (12/01/10).
Preparing for NEODC.

comment:27 Changed 9 years ago by knpa

Tarball created and NEODC informed ready for transfer.

comment:28 Changed 9 years ago by knpa

Transferred to NEODC.
Closing ticket.

comment:29 Changed 9 years ago by knpa

  • Resolution set to fixed
  • Status changed from assigned to closed
Note: See TracTickets for help on using tickets.